### v1

@@ Abstract
Anatomical changes in head-and-neck cancer (HNC) patients during intensity-modulated proton therapy (IMPT) can shift the Bragg Peak of proton beams, risking tumor underdosing and organ-at-risk (OAR) overdosing. As a result, treatment replanning is often required to maintain clinically acceptable treatment quality. However, current manual replanning processes are often resource-intensive and time-consuming. In this work, we propose a patient-specific deep reinforcement learning (DRL) framework for automated IMPT replanning, with a reward-shaping mechanism based on a -point plan quality score designed to handle the competing clinical objectives in radiotherapy planning. We formulate the planning process as an RL problem where agents learn high-dimensional control policies to adjust plan optimization priorities to maximize plan quality. Unlike population-based approaches, our framework trains personalized agents for each patient using their planning CT (Computed Tomography) and augmented anatomies simulating anatomical changes (tumor progression and regression). This patient-specific approach leverages anatomical similarities along the treatment course, enabling effective plan adaptation. We implemented and compared two DRL algorithms, Deep Q-Network (DQN) and Proximal Policy Optimization (PPO), using dose-volume histograms (DVHs) as state representations and a -dimensional action space of priority adjustments. Evaluation on five HNC patients using actual replanning CT data showed that both DRL agents improved initial plan scores from to (DQN) and (PPO), surpassing the replans manually generated by a human planner (). Clinical validation confirms these improvements translate to better tumor coverage and OAR sparing across diverse anatomical changes. This work highlights the potential of DRL in addressing the geometric and dosimetric complexities of adaptive proton therapy, offering a promising solution for efficient offline adaptation and paving the way for online adaptive proton therapy.
@@ 1 Introduction
Intensity-modulated proton therapy (IMPT) provides highly conformal tumor coverage while sparing surrounding organs at risk (OARs), leveraging the unique dose deposition characteristics of the proton beam’s Bragg peak (Holliday et al., 2015; McKeever et al., 2016; Moreno et al., 2019). However, the Bragg peak makes IMPT highly sensitive to interfractional anatomical changes, such as tumor progression or regression, weight loss, and edema, that alter tissue density along the beam path (Huiskes et al., 2023; Sonke et al., 2019). These anatomical changes can shift the Bragg peak, degrading tumor coverage and/or OAR sparing, and often necessitate one or multiple manual replanning sessions during treatment to maintain clinical quality. Addressing these anatomical changes currently requires manual intervention: a multi-step process involving patient reimaging, anatomical recontouring, and iterative, trial-and-error plan optimization. This manual adaptation workflow is notoriously labor-intensive and typically takes several days, delaying optimal care and straining limited clinical resources (Kim et al., 2018; Leeman et al., 2017; Li et al., 2020).
@@
Reinforcement learning (RL) offers a compelling framework for tackling these sequential, multi-objective optimization challenges in HNC treatment planning. RL agents learn optimal strategies through interaction and reward feedback, making them particularly well-suited for navigating the complex, iterative process of balancing competing clinical objectives in treatment planning (Sutton et al., 1998). Specifically, RL holds the potential to automate the iterative decision-making process of priority adjustments in treatment planning (Eckardt et al., 2021; Ebrahimi and Lim, 2021; Moreau et al., 2021; Yang et al., 2024). Furthermore, RL’s inherent adaptability aligns with the vision of personalized online adaptive radiotherapy, where treatment plans can be dynamically adapted based on a patient’s tumor response and anatomical changes during treatment.
To realize this potential for automated IMPT replanning, we develop a patient-specific deep reinforcement learning (DRL) framework for the dosimetrically challenging HNC. Our approach formulates the priority tuning process during plan optimization as an RL problem, where the state is defined by dose–volume histograms (DVHs) for clinical target volumes (CTVs) and OARs, and the action space comprises predefined, clinically informed priority adjustments, enabling the agent to navigate the trade-offs inherent in this multi-objective optimization problem. Plan quality is quantified via a comprehensive -point scoring system that combines ProKnow standardized scoring criteria (Nelms et al., 2012) with institutional planning guidelines, and the reward is defined as the change in the plan quality score. Our framework trains personalized DRL agents using each patient’s anatomy captured by the initial planning CT (Computed Tomography) images and augmented anatomies simulating anatomical variations (tumor regression/progression). This patient-specific approach leverages the patient’s inherent anatomical consistency along the treatment course to optimize performance specifically for that individual rather than attempting to develop a one-size-fits-all solution.
@@ 2 Related Work
HNC represents a substantially more challenging planning context due to multiple target volumes with varying prescription doses and numerous critical structures in close proximity, demanding tight dosimetric trade-offs even for experienced planners. Recent RL adaptations to HNC planning have attempted to navigate this high-dimensional optimization space with varied strategies. Gao et al. (2024) adapted the hierarchical VTP approach for HNC photon therapy (whereas this work focuses on proton therapy), employing two DQN subnetworks for parameter selection and adjustment direction determination while representing states as plan quality scores rather than DVHs. To manage the high dimensionality arising in HNC planning, they manually fixed most of the potential planning parameters in their case based on dosimetrist expertise, allowing their VTP to learn adjustments for only unique priority parameters. Similarly, grappling with the high-dimensional action space, Wang and Chang (2024) explored policy-gradient methods with a transformer-based PPO agent for continuous parameter adjustments. Their approach proposed dynamically reducing complexity by adjusting parameters only for the subset of structures currently exhibiting the lowest quality scores. However, this strategy may not accommodate cases where certain OARs must be strategically compromised to maintain target coverage—a common requirement in complex HNC planning scenarios.
Notably, all the aforementioned approaches rely on population-based models trained across diverse patient cohorts. While effective at capturing common anatomical patterns, such models face challenges adapting to patient-specific anatomical changes that necessitate treatment replanning during HNC’s fractionated radiotherapy (Caudell et al., 2017; Ma et al., 2022; Maniscalco et al., 2023; Visak et al., 2024). This limitation, stemming from the under-utilization of intra-patient anatomical consistency, restricts their ability to generate optimally adapted plans tailored to each patient’s unique progression. Motivated by these challenges, we introduce a novel patient-specific RL framework for IMPT replanning in HNC. By developing and optimizing the planning policy to each individual’s unique anatomy and dosimetric constraints from the outset, our approach inherently leverages intra-patient consistency patterns. This personalization is crucial for effectively managing the high-dimensional optimization space and is potentially better suited to addressing the adaptive replanning demands inherent in HNC treatment, overcoming key limitations of existing population-based strategies.
@@ 3.1 Treatment Plan Optimization
Here, is the multi-objective treatment planning loss function, where is the vector of beamlet intensities to be optimized. Each component represents the intensity or weight of the -th beamlet. The constraint explicitly ensures that all beamlet intensities are non-negative. The formulation incorporates CTVs, where indexes each CTV. Similarly, OARs are considered, where indexes each OAR. For the patient cohort in this study, and . The weighting parameters and serve as treatment planning priorities (TPPs) that balance competing clinical objectives: achieving prescribed dose coverage in the -th CTV while minimizing dose exposure to the -th OAR. These weights are the primary optimization variables manipulated by our DRL agents. and denote the dose influence matrices for the -th CTV and -th OAR. is the prescribed dose vector for the -th CTV.
In practice, achieving a perfect homogeneous dose distribution within CTVs is not feasible; hot and cold spots are often unavoidable. Similarly, since radiation beams must often pass through some OARs to reach the CTVs, complete OAR sparing from radiation dose is unrealistic. Importantly, different OARs have varying dose tolerances based on their biological characteristics; as long as the dose to each OAR remains within clinically acceptable thresholds, the risk of severe toxicity is minimized. Hence, to reflect these clinical realities and achieve realistic plan optimization, in this study we implemented a DVH-based cost function incorporating multiple planning objectives. The planning objectives involved in our IMPT plan optimization for HNC are listed in Table 1, where represents the volume of a structure receiving a dose equal to or greater than , while denotes the minimum dose received by the hottest volume of a structure. The trade-offs between these competing objectives are controlled by objective-specific priorities (), which are traditionally set manually through iterative tuning. The DVH-based cost function (Eqn. 2) was optimized using an in-house dose calculation engine, which serves as the environment for our RL-based planning optimization.
@@ 3.2 Reinforcement Learning for Automatic Replanning
Action Space (): A discrete set of clinically constrained actions representing priority adjustments for different planning objectives, designed based on clinical experience.
@@
Reward Function (): The reward function defines the immediate reward obtained after taking action in state . It is defined as , where is the plan quality score of the intermediate plan dose, generated by taking action . We designed a -point scoring system based on the standardized ProKnow scoring criteria (Nelms et al., 2012) and institution-specific planning guidelines. It includes a set of dosimetric metrics for the structures of interest. The final score was calculated as the sum of the scores across all these metrics, with higher scores indicating better plan quality.
@@
Proximal Policy Optimization (PPO)
@@ 4.1 Cohort Selection
We validated our approach through a retrospective study involving five HNC patients treated with IMPT at Hospital A. All patients underwent replanning during their treatment course due to substantial anatomical changes. A CT scan for treatment planning (referred to as pCT) was acquired weeks before the treatment course, and a second CT scan was obtained for replanning (rpCT) during the course when substantial anatomical changes were observed and warranted replanning. These five cases were deliberately selected for their varied tumor characteristics and spatial relationships to surrounding OARs, and all cases involved three CTVs, representing challenging scenarios for treatment planning. The primary CTV (CTV1) for all cases was prescribed to receive over fractions, while the prescription doses for the secondary (CTV2) and tertiary (CTV3) targets varied among cases, as detailed in Table 2. This variability in prescription doses poses a significant challenge for automated treatment planning and replanning in HNC, as population-based models often struggle to generalize across diverse clinical objectives, let alone the substantial variation in patient anatomy and tumor characteristics.
Case ID | Volume on pCT () | Volume on rpCT () | No. of Fractions | Prescription (Gy RBE) | ||||||
CTV1 | CTV2 | CTV3 | CTV1 | CTV2 | CTV3 | CTV1 | CTV2 | CTV3 | ||
P1 | 35 | |||||||||
P2 | 35 | |||||||||
P3 | 35 | |||||||||
P4 | 35 | |||||||||
P5 | 35 |
@@ 4.2 Data Extraction
To train a patient-specific agent capable of adjusting the priorities among planning objectives to adapt the treatment plan to anatomical changes, we further generated augmented anatomies by simulating anatomical changes. In this proof-of-concept study, we simulated only two variation scenarios, i.e., tumor progression and regression, by expanding the original CTV contours on the pCT by or shrinking them by , respectively. These variations not only changed the CTV volumes but also altered their spatial relationships with surrounding OARs, increasing the diversity of the patient-specific training dataset in terms of anatomy and planning complexity. The pCT images, along with each set of modified contours, were also imported into matRad to calculate the corresponding dose influence matrices for the simulated anatomy variations.
@@ 5 Experiments & Results
Case ID | Initial plan | Manual plan | DQN plan | PPO plan |
P1 | 122.56 | 130.85 | 131.05 | 137.88 |
P2 | 83.14 | 132.09 | 136.20 | 138.05 |
P3 | 132.35 | 140.44 | 147.25 | 147.52 |
P4 | 132.14 | 143.94 | 146.10 | 148.77 |
P5 | 132.96 | 138.67 | 138.32 | 141.50 |
@@
5.1 Experimental Setup
For each patient, both DQN and PPO agents were trained using the patient’s original anatomy from the pCT and two augmented anatomical variations. The state representation included the normalized DVH curves for all structures (dim=), while the discrete action space consisted of priority adjustments for CTVs and OARs. Each agent was trained for episodes, with each episode consisting of up to priority adjustment steps or terminating early if the maximum plan quality score of was achieved. Algorithm 1 provides a detailed overview of this DRL-based replanning process, illustrating the integration of priority tuning and plan optimization within our experimental framework.
The trained agents were then tested on the patient’s new anatomy captured in the rpCT. Specifically, starting from an initial default priority set , the agents performed priority tuning within tuning steps, where each step involved selecting a tuning action based on the plan quality of the intermediate plan and was followed by inverse plan optimization. For comparison, a new treatment plan was also manually generated for each case on the new anatomy by a human planner through manual priority tuning. The performance of the agents was assessed by comparing the plan quality scores and dosimetric metrics of the agent-generated plans with those of the manually created plans.
To facilitate the interaction between the RL agents and the treatment planning process, we developed a custom OpenAI Gymnasium-compatible environment that simulates the treatment planning workflow. At each timestep, the environment processes the current beamlet weights along with the dose influence matrices and for each CTV and OAR . It then solves the DVH-based optimization problem using the priority set , modified by the agent-selected action , updates , and generates a new dose distribution , which is subsequently used to compute the plan score and reward via Eqn. 1. Each action updates priority weights according to:
where ensures weights remain within valid bounds. All experiments were conducted on a workstation equipped with an NVIDIA RTX Ada Generation GPU with of memory.

### v2

@@ Abstract
Anatomical changes in head-and-neck cancer (HNC) patients during intensity-modulated proton therapy (IMPT) can shift the Bragg Peak of proton beams, risking tumor underdosing and organ-at-risk (OAR) overdosing. As a result, treatment replanning is often required to maintain clinically acceptable treatment quality. However, current manual replanning processes are often resource intensive and time consuming. In this work, we propose a patient-specific deep reinforcement learning (DRL) framework for automated IMPT replanning, with a reward-shaping mechanism based on a -point plan quality score designed to handle competing clinical objectives in radiotherapy planning. We formulate the planning process as a reinforcement learning (RL) problem where agents learn high-dimensional control policies to adjust plan optimization priorities to maximize plan quality. Unlike population-based approaches, our framework trains personalized agents for each patient using their planning Computed Tomography (CT) and augmented anatomies simulating anatomical changes (tumor progression and regression). This patient-specific approach leverages anatomical similarities along the treatment course, enabling effective plan adaptation. We implemented and compared two DRL algorithms, Deep Q-Network (DQN) and Proximal Policy Optimization (PPO), using dose-volume histograms (DVHs) as state representations and a -dimensional action space of priority adjustments. Evaluation on eight HNC patients using actual replanning CT data showed that both DRL agents improved initial plan scores from to (DQN) and (PPO), surpassing the replans manually generated by a human planner (). Further comparison of dosimetric endpoints confirms these improvements translate to better tumor coverage and OAR sparing across diverse anatomical changes. This work highlights the potential of DRL in addressing the geometric and dosimetric complexities of adaptive proton therapy, offering a promising solution for efficient offline adaptation and paving the way for online adaptive proton therapy.
@@ 1 Introduction
Intensity-modulated proton therapy (IMPT) provides highly conformal tumor coverage while sparing surrounding organs at risk (OARs), leveraging the unique dose deposition characteristics of the proton beam’s Bragg peak (Holliday et al., 2015; McKeever et al., 2016; Moreno et al., 2019). However, the Bragg peak makes IMPT highly sensitive to interfractional anatomical changes, such as tumor progression or regression, weight loss, and edema, that alter tissue density along the beam path (Huiskes et al., 2023; Sonke et al., 2019). These anatomical changes can shift the Bragg peak, degrading tumor coverage and/or OAR sparing, and often necessitate one or multiple manual replanning sessions during the treatment course to maintain clinical quality. Addressing these anatomical changes currently requires manual intervention: a multi-step process involving patient reimaging, anatomical recontouring, and iterative, trial-and-error plan optimization. This manual adaptation workflow is notoriously labor intensive and typically takes several days, delaying optimal care and straining limited clinical resources (Kim et al., 2018; Leeman et al., 2017; Li et al., 2020).
@@
Reinforcement learning (RL) offers a compelling framework for tackling these sequential, multi-objective optimization challenges in HNC treatment planning. RL agents learn optimal strategies through interaction and reward feedback, making them particularly well-suited for navigating the complex, iterative process of balancing competing clinical objectives in treatment planning (Sutton et al., 1998). Specifically, RL holds the potential to automate the iterative decision-making process of priority adjustments in treatment planning (Eckardt et al., 2021; Ebrahimi and Lim, 2021; Moreau et al., 2021; Yang et al., 2024). Furthermore, RL’s inherent adaptability aligns with the vision of personalized online adaptive radiotherapy, where treatment plans can be dynamically adapted based on a patient’s tumor response and anatomical changes along the treatment course.
To realize this potential for automated IMPT replanning, we develop a patient-specific deep reinforcement learning (DRL) framework for the dosimetrically challenging HNC. Our approach formulates the priority tuning process during plan optimization as an RL problem, where the state is defined by dose–volume histograms (DVHs) for clinical target volumes (CTVs) and OARs, and the action space comprises predefined, clinically informed priority adjustments, enabling the agent to navigate the trade-offs inherent in this multi-objective optimization problem. Plan quality is quantified via a comprehensive -point scoring system that combines ProKnow standardized scoring criteria (Nelms et al., 2012) with institutional planning guidelines, and the reward is defined as the change in the plan quality score. Our framework trains personalized DRL agents using each patient’s anatomy captured by the initial planning Computed Tomography (CT) images and augmented anatomies simulating anatomical variations. This patient-specific approach leverages the patient’s inherent anatomical consistency along the treatment course to optimize performance specifically for that individual rather than attempting to develop a one-size-fits-all solution.
@@ 2 Related Work
HNC represents a substantially more challenging planning context due to multiple target volumes with varying prescription doses and numerous critical structures in close proximity, demanding tight dosimetric trade-offs even for experienced planners. Recent RL adaptations to HNC planning have attempted to navigate this high-dimensional optimization space with varied strategies. Gao et al. (2024) adapted the hierarchical VTP approach for HNC photon therapy (whereas our work focuses on proton therapy), employing two DQN subnetworks for parameter selection and adjustment direction determination while representing states as plan quality scores rather than DVHs. To manage the high dimensionality arising in HNC planning, they manually fixed most of the potential planning parameters in their case based on dosimetrist expertise, allowing their VTP to learn adjustments for only unique priority parameters. Similarly, grappling with the high-dimensional action space, Wang and Chang (2024) explored policy-gradient methods with a transformer-based PPO agent for continuous parameter adjustments. Their approach proposed dynamically reducing complexity by adjusting parameters only for the subset of structures currently exhibiting the lowest quality scores. However, this strategy may not accommodate cases where certain OARs have to be strategically compromised to maintain target coverage—a common situation in complex HNC planning scenarios.
Notably, all the aforementioned approaches rely on population-based models trained across diverse patient cohorts. While effective at capturing common anatomical patterns, such models face challenges adapting to patient-specific anatomical changes that necessitate treatment replanning during HNC’s fractionated radiotherapy (Caudell et al., 2017; Ma et al., 2022; Maniscalco et al., 2023; Visak et al., 2024). This limitation, stemming from the under-utilization of intra-patient anatomical consistency, restricts their ability to generate optimally adapted plans tailored to each patient’s unique anatomical features and evolving pattern. Motivated by these challenges, we introduce a novel patient-specific RL framework for IMPT replanning in HNC. By developing and optimizing the planning policy to each individual’s unique anatomy and dosimetric constraints from the outset, our approach inherently leverages intra-patient consistency patterns. This personalization is crucial for effectively managing the high-dimensional optimization space and is potentially better suited to addressing the adaptive replanning demands inherent in HNC treatment, overcoming key limitations of existing population-based strategies.
@@ 3.1 Treatment Plan Optimization
Here, is the multi-objective treatment planning loss function, where is the vector of beamlet intensities to be optimized. Each component represents the intensity or weight of the -th beamlet. The constraint explicitly ensures that all beamlet intensities are non-negative. The formulation incorporates CTVs, where indexes each CTV. Similarly, OARs are considered, where indexes each OAR. For the patient cohort in this study, and . and denote the dose influence matrices for the -th CTV and -th OAR. is the prescribed dose vector for the -th CTV.
The weighting parameters and serve as treatment planning priorities (TPPs) that balance competing clinical objectives: achieving prescribed dose coverage in the -th CTV while minimizing dose exposure to the -th OAR. While this weighted loss function is a standard proxy for the clinical objective, it does not directly optimize for the clinical constraints that human planners must satisfy (e.g., specific DVH criteria for various organs listed in Table 1). In practice, achieving perfect dose homogeneity is not feasible, and different OARs have varying dose tolerances based on their biological characteristics. These clinical realities mean that planners must manually and iteratively adjust the weighting parameters and re-solve the optimization problem until clinical requirements are met–a time-consuming and resource-intensive process.
We automate this critical step by introducing a DRL agent to perform this hyperparameter search. The agent learns to directly optimize for clinical constraint satisfaction, using a reward signal derived from the dose plan’s adherence to clinical goals. The DRL agent interacts with a plan optimization engine that serves as the core of our RL environment, minimizing the cost function (Eqn. 2) with agent-selected priority weights. This integration of dose calculation with RL-based planning optimization streamlines a major bottleneck in the radiotherapy workflow.
@@ 3.2 Reinforcement Learning for Automatic Replanning
Action Space (): A discrete set of clinically constrained actions representing priority adjustments for different planning objectives, designed based on clinical experience. Adjustment details are provided in Appendix C.
@@
Reward Function (): The reward function defines the immediate reward obtained after taking action in state . It is defined as , where is the plan quality score of the intermediate plan dose, generated by taking action . We designed a -point scoring system based on the standardized ProKnow scoring criteria (Nelms et al., 2012) and institution-specific planning guidelines. It includes a set of dosimetric metrics for the structures of interest. The final score was calculated as the sum of the scores across all these metrics, with higher scores indicating better plan quality. The plot of the scoring function for each dosimetric metric is presented in Appendix D.
@@
Proximal Policy Optimization (PPO):
@@ 4.1 Cohort Selection
We validated our approach using a retrospective cohort of HNC patients treated with IMPT at the Emory Proton Therapy Center. Eight patients requiring replanning due to substantial anatomical changes were used for patient-specific RL training, while the remaining 10 patients served as training data for population-based RL baselines. For each patient, an initial planning CT (pCT) was acquired weeks before the treatment course, and a replanning CT (rpCT) was obtained when significant anatomical changes were observed and warranted replanning. Performance evaluation used all eight rpCT cases for both RL approaches. These cases were selected for their varied tumor characteristics and spatial relationships to surrounding OARs, all involving three CTVs representing challenging planning scenarios. While the primary CTV (CTV1) for all cases was prescribed over fractions, the prescription doses for the secondary (CTV2) and tertiary (CTV3) targets varied per case (Table 2). This prescription heterogeneity, combined with substantial inter-patient anatomical variation, poses a major challenge for automated treatment planning in HNC, particularly for population-based models struggling to generalize across heterogeneous clinical objectives. Patient demographics and staging are detailed in Table 6 (Appendix B).
[…]
CTV1 | CTV2 | CTV3 | CTV1 | CTV2 | CTV3 | CTV1 | CTV2 | CTV3 | CTV1 | CTV2 | CTV3 | ||
P1 | 2 | ||||||||||||
P2 | 1 | NA | NA | NA | |||||||||
P3 | 1 | NA | NA | NA | |||||||||
P4 | 2 | ||||||||||||
P5 | 1 | NA | NA | NA | |||||||||
P6 | 2 | ||||||||||||
P7 | 2 | ||||||||||||
P8 | 2 |
@@ 4.2 Data Extraction
To train a patient-specific agent capable of adapting treatment plans to anatomical changes, we generated augmented anatomies by simulating tumor variations. In this proof-of-concept study, we simulated only two variation scenarios, i.e., tumor progression and regression, by expanding the original CTV contours on the pCT by or shrinking them by , respectively. These variations not only changed the CTV volumes but also altered their spatial relationships with surrounding OARs, increasing the diversity of the patient-specific training dataset in terms of anatomy and planning complexity. The pCT images, along with each set of modified contours, were also imported into matRad to calculate the corresponding dose influence matrices for the simulated anatomy variations.
@@ 5 Experiments & Results
5.1 Experimental Setup
For each patient, both DQN and PPO agents were trained using the patient’s original anatomy from the pCT and two augmented anatomical variations. The state representation included the normalized DVH curves for all structures (dim=), while the discrete action space consisted of priority adjustments for CTVs and OARs. Each agent was trained for episodes, with each episode consisting of up to priority adjustment steps or terminating early if the maximum plan quality score of was achieved. Algorithm 1 in Appendix A provides a detailed overview of this DRL-based replanning process, illustrating the integration of priority tuning and plan optimization within our experimental framework. Details of network specifications and hyperparameters are provided in Appendix E.
The trained agents were then tested on the patient’s new anatomy captured in the rpCT. Specifically, starting from an initial default priority set , the agents performed priority tuning following the same episodic framework described above, where each step involved selecting a tuning action based on the intermediate plan’s quality, followed by inverse plan optimization. The number of tuning attempts (horizon length) in this process can be extended as needed, bounded only by computational cost. As a comparative baseline, a human planner also generated a new manual treatment plan for each case, employing manual priority tuning. The performance of the agents was assessed by comparing the plan quality scores and dosimetric metrics of the agent-generated plans with those of the manually created plans.
To facilitate the interaction between the RL agents and the treatment planning process, we developed a custom OpenAI Gymnasium-compatible environment that simulates the treatment planning workflow. At each time step, the environment receives the agent-selected action at and applies it to modify the current priority weights . Using these updated priorities along with the dose influence matrices and for each CTV and OAR , the environment solves the optimization problem (Eqn. 2) via projected gradient descent (Fu et al., 2023; Ghobadi et al., 2012; Nocedal and Wright, 1999) to update the beamlet weights and generate a new dose distribution (Eqn. 1). The resulting dose distribution is evaluated against clinical constraints (Table 1) to compute the plan score and reward (Appendix D). Each action updates priority weights according to
where ensures weights remain within valid bounds. The details of the priority adjustment are included in the Appendix C. All experiments were conducted on a workstation equipped with an NVIDIA RTX Ada Generation GPU with of memory.
Case ID | Initial | Manual | DQN-popn | DQN | PPO-popn | PPO |
P1 | 122.56 | 130.85 | 130.40 | 131.05 | 137.01 | 137.88 |
P2 | 83.14 | 132.09 | 135.33 | 136.20 | 135.64 | 138.05 |
P3 | 132.35 | 140.44 | 144.17 | 147.25 | 147.31 | 147.52 |
P4 | 132.14 | 143.94 | 148.38 | 146.10 | 146.38 | 148.77 |
P5 | 132.96 | 138.67 | 138.41 | 138.32 | 138.30 | 141.50 |
P6 | 125.65 | 133.03 | 128.87 | 135.24 | 135.18 | 135.18 |
P7 | 108.67 | 132.53 | 126.27 | 141.03 | 136.51 | 141.35 |
P8 | 128.75 | 139.03 | 124.31 | 141.53 | 137.06 | 141.78 |